Pamela requires 3 hours to swim 15 miles downstream in the illinois river. The return trip upstream takes 5 hours. Find Pamela's average speed in still water. How fast is the current?( assume that Pamela's speed is the same in each direction)
Answer by stanbon(57984)

Pamela requires 3 hours to swim 15 miles downstream in the illinois river.
-----
time = 3 hrs; distance = 15 miles ; rate = d/t = 15/3 = 5 mph
-----------------------------
The return trip upstream takes 5 hours.
time = 5 hrs; distance = 15 miles ; rate = 15/5 = 3 mph
------------------------------
Find Pamela's average speed in still water.
How fast is the current?( assume that Pamela's speed is the same in each direction)
---
Let her swim speed be "s":
Let the current speed be "c";
Equations:
Down stream: s + c = 5 mph
up-stream::: s - c = 3 mph
-----------------
Add the two equations and solve for "s"
2s = 8 mph
s = 4 mph (Pamel's speed in still water.
---
Since s+c=5, c = 1 mph (speed of the current)
=====================
Cheers,
Stan H.
=====================